TORSIONAL ANALYSIS OF 2-BUTYNOL AND 2-PENTYNE

\underline{RANGA SUBRAMANIAN}, Department of Chemistry, Wesleyan University, Middletown, CT 06459; ROBERT K. BOHN, Department of Chemistry, University of Connecticut, Storrs, CT 06269 ; and STEWART E. NOVICK, Department of Chemistry, Wesleyan University, Middletown, CT 06459.

The microwave spectrum of 2-butynol, CH_{3}-C≡C-CH_{2}OH, has been investigated in the frequency range of 6 to 26.5 GHz. The spectra of the A and E torsional states were observed using Fourier transform microwave spectroscopy. Due to the presence of the cylindrically symmetric -C≡C- ”spacer” between the methyl group and the rest of the molecule, the barrier to internal rotation, V_{3}, is only 7 cm^{−1}. One conformer of 2-butynol was observed and assigned. The spectrum was analyzed with the ρ-axis method using a very flexible Hamiltonian which gives a fit an order of magnitude better than that obtained with more standard code for internal rotation. The spectroscopic constants for 2-butynol are 23744.(18), 2093.429(1), 1966.358(1), and -400.34(2) MHz for A, B, C, and D_{ab} ; 0.48(1) kHz, -30.3(4) kHz, and 4.5(5) MHz for Δ_{J}, Δ_{JK}, and Δ_{K}; and 0.1406(6), 6.93(9) cm^{−1}, -33.4(7) kHz, and 192.0(9) GHz for ρ, V_{3}, L_{V}, and F, respectively. The root mean squared error of the fit is 14 kHz. The angles between the internal rotor axis and the principal axes are θ_{a} = 12^{∘}, θ_{b} = 78^{∘}, with θ_{c} held fixed at 90^{∘}. A preliminary fit for the torsional states of 2-pentyne, CH_{3}-C≡C-CH_{2}CH_{3}, will also be presented.